# Towards Enhanced Immersion and Agency for LLM-based Interactive Drama

### v1

@@ Abstract
LLM-based Interactive Drama is a novel AI-based dialogue scenario, where the user (i.e. the player) plays the role of a character in the story, has conversations with characters played by LLM agents, and experiences an unfolding story. This paper begins with understanding interactive drama from two aspects: Immersion—the player’s feeling of being present in the story—and Agency—the player’s ability to influence the story world. Both are crucial to creating an enjoyable interactive experience, while they have been underexplored in previous work. To enhance these two aspects, we first propose Playwriting-guided Generation, a novel method that helps LLMs craft dramatic stories with substantially improved structures and narrative quality. Additionally, we introduce Plot-based Reflection for LLM agents to refine their reactions to align with the player’s intentions. Our evaluation relies on human judgment to assess the gains of our methods in terms of immersion and agency.111github.com/gingasan/interactive-drama
@@
Hongqiu Wu, Weiqi Wu, Tianyang Xu, Jiameng Zhang, Hai Zhao Department of Computer Science and Engineering, Shanghai Jiao Tong University wuhongqiu@sjtu.edu.cn
@@ 1 Introduction
The above theory provides foundation for our work. We introduce the immersion-agency paradigm for LLM-based interactive drama in Figure 2. To enhance immersion and agency based on this paradigm, we make two key contributions.
@@ 2 Related Work
It has been widely used in recent years to scale the training performance by using AI-generated data (Wang et al., 2023b; Lee et al., 2023; Wu et al., 2024). Our research focuses on the use of LLMs to assist humans in creating high-quality literature, such as poetry (Ormazabal et al., 2022; Chakrabarty et al., 2022), music (Hernandez-Olivan et al., 2022; Zhu et al., 2023), stories (Tan et al., 2021; Tian et al., 2024). Our method generates a story given a short paragraph. This is more challenging than generating the drama script given a complete story (Zhao et al., 2024).
@@ 4.1 Preliminary
Let us regard the entirety of role agents as a black-box decision function ${\rm Decision}$. At each moment $t$, it processes the memories, etc., takes the player input $x_{t}$, and makes a decision (e.g. speak something to someone) $y_{t}={\rm Decision}(x_{t},m_{t},o_{t},\sigma)$, where $m_{t}$, $o_{t}$, and $\sigma$ denote the memories, observations, and drama script, respectively.
@@ 4.3 Architecture
Figure 3: Hybrid architecture and plot-based reflection. The first scene emphasizes background telling. We apply the one-for-all architecture. The following scenes involve mutual suspicion and personal investigation. We apply the director-actor architecture.
@@
To balance performance and efficiency, we apply a hybrid architecture, which dynamically switches between two architectures based on the characteristic of the scene, as shown in Figure 3. For scenes that emphasize interaction (e.g. exploring clues), the director-actor architecture remains the optimal choice. For scenes that emphasize narrative, we simplify it to one-for-all, which means the director agent reflects and makes the decision directly.
@@ 4.4 Other Implementation Details
There are other important details to implement the role agents. $\bullet$ Model: As in previous work (Park et al., 2023; Zhao et al., 2024), we prompt the state-of-the-art LLM to construct agents. In this work, the director and actor agents are all based on GPT-4o. It is a general LLM instead of reasoning LLM like OpenAI-o1 and DeepSeek-R1 (DeepSeek-AI, 2025). $\bullet$ Memory: We summarize the memories of each agent at the end of each scene. During the scene, we retain all memories in the prompt. $\bullet$ Transition: A scene will transit to the next when all plots in the plot chain have been completed.
@@
Playwriting-Guided 32% 24% 32% 22% 48% 16% 34% 20% 66%
    w/o. critic & revise 24% 24% 26% 34% 18% 28% 12% 32% 92%
@@
Table 3: Win-rates of different dramatic story generation methods over 50 topics. “B” and “W” refer to the best and the worst. We highlight the best and worst score using bold and underline respectively for each dimension.
@@ 5.2 Evaluation on Role Agents
Table 4 compares the performance of different architectures. For each architecture, we let each human/agent play the drama once. This results in 60 playing histories totally. We report the average score of each dimension. The hybrid architecture refers to applying one-for-all for the first scene and director-actor for the second and third scene. We first focus on the effect of plot-based reflection. It significantly enhances agency from both dimensions (3.5$\rightarrow$4.0 and 3.3$\rightarrow$4.0). We also notice that plot-based reflection is useful in enhancing character attractiveness (3.5$\rightarrow$3.9) and narrative progression (3.9$\rightarrow$4.3). We conjecture that the reflection encourages characters to display a stronger sense of empathy, which thus contributes to higher attractiveness and decent progression.
The last column of Table 4 suggests the efficiency boost achieved by the hybrid architecture. We find that the director-character architecture is advantageous in building character attractiveness because the independent thinking process is more likely to give impressive responses. However, it incurs a lower quality of narrative, especially in progression (4.3$\rightarrow$3.6). This is due to the fact that the communication between director and actor agents inevitably results in information loss.
@@
Table 5 further indicates that the hybrid architecture remains decent in the face of two types of players. Interestingly, it even higher scores in character attractiveness and player influence against aggressive agents compared to human players. We conjecture that this is because agents are more eager to interact and the high-quality responses in turn impress the annotators. However, the provocative queries might lead to missing of plots, as shown by the drop in narrative progression.
@@ 6 Conclusion
Our paper adopts a limited scope of playwriting techniques. However, in modern drama, they should not be predetermined; they should evolve and be adapted based on various domains. There is a potential gap between two parts of evaluation. We evaluate dramatic story generation based on 50 manually-crafted premises, while role agents are evaluated based on a high-quality script. We have not found an efficient method to quantify the player’s complete experience across both the creation and engagement of the story. For example, inviting the volunteers to craft the story, play the interactive drama, and eventually make comments can be inefficient and unstable. For our demonstration, we develop visual scenery using the Phaser3 game engine. However, the drama we study in this paper focuses on text-based interactive experiences. Our method is based on prompting GPT-4o, one of the most powerful LLMs, but it still exhibits shortcomings in plot-based reflection and making ideal character decisions. Our future work will focus on developing effective data-driven methods to enhance the specific capabilities. Our experiments are based on GPT-4o, without extending to a broader range of popular LLMs, e.g. Claude, DeepSeek.

### v2

@@ Abstract
LLM-based Interactive Drama is a novel AI-based dialogue scenario, where the user (i.e. the player) plays the role of a character in the story, has conversations with characters played by LLM agents, and experiences an unfolding story. This paper begins with understanding interactive drama from two aspects: Immersion—the player’s feeling of being present in the story—and Agency—the player’s ability to influence the story world. Both are crucial to creating an enjoyable interactive experience, while they have been underexplored in previous work. To enhance these two aspects, we first propose Playwriting-guided Generation, a novel method that helps LLMs craft dramatic stories with substantially improved structures and narrative quality. Additionally, we introduce Plot-based Reflection for LLM agents to refine their reactions to align with the player’s intentions. Our evaluation relies on human judgment to assess the gains of our methods in terms of immersion and agency.
@@
Hongqiu Wu1,3,4, Weiqi Wu1,3,4, Tianyang Xu1,3,4, Jiameng Zhang2, Hai Zhao1,3,4 1School of Computer Science, Shanghai Jiao Tong University 2University of Zurich 3 Key Laboratory of Shanghai Education Commission for Intelligent Interaction and Cognitive Engineering, Shanghai Jiao Tong University 4 Shanghai Key Laboratory of Trusted Data Circulation and Governance in Web3 {wuhongqiu,wuwq1022,johnnie.walker}@sjtu.edu.cn, jiameng.zhang@uzh.ch,zhaohai@cs.sjtu.edu.cn
Corresponding author. This research was supported by the Joint Research Project of Yangtze River Delta Science and Technology Innovation Community (No. 2022CSJGG1400). Code and demonstration are in https://github.com/gingasan/interactive-drama.
@@ 1 Introduction
The above theory provides a foundation for our work. We introduce the immersion-agency paradigm for LLM-based interactive drama in Figure 2. To enhance immersion and agency based on this paradigm, we make two key contributions.
@@ 2 Related Work
It has been widely used in recent years to scale the training performance by using AI-generated data (Wang et al., 2023b; Lee et al., 2023; Wu et al., 2024). Our research focuses on the use of LLMs to assist humans in creating high-quality literature, such as poetry (Ormazabal et al., 2022; Chakrabarty et al., 2022), music (Hernandez-Olivan et al., 2022; Zhu et al., 2023), stories (Tan et al., 2021; Tian et al., 2024; Riedl, 2016). It is related to computational creativity (Wiggins, 2006; Gervás, 2009; Colton and Wiggins, 2012). Our method generates a story given a short paragraph. This is more challenging than generating the drama script given a complete story (Zhao et al., 2024).
@@ 4.1 Preliminary
Let us regard the entirety of role agents as a black-box decision function ${\rm Decision}$. At each moment $t$, it processes some text inputs and makes a decision (e.g. speak something to someone) $y_{t}={\rm Decision}(x_{t},m_{t},o_{t},\sigma)$, where $x_{t}$, $m_{t}$, $o_{t}$, and $\sigma$ denote the player input, memories, observations, and drama script, respectively.
@@ 4.3 Architecture
Figure 3: and plot-based reflection. The first scene emphasizes background telling. We apply the one-for-all architecture. The following scenes involve mutual suspicion and personal investigation. We apply the director-actor architecture.
@@
To balance performance and efficiency, we apply a , which dynamically switches between two architectures based on the characteristic of the scene, as shown in Figure 3. For scenes that emphasize interaction (e.g. exploring clues), the director-actor architecture remains the optimal choice. For scenes that emphasize narrative, we simplify it to one-for-all, which means the director agent reflects and makes the decision directly.
@@ 4.4 Other Implementation Details
There are other important details to implement the role agents. $\bullet$ Model: As in previous work (Park et al., 2023; Zhao et al., 2024), we prompt the state-of-the-art LLM to construct agents. In this work, the director and actor agents are all based on GPT-4o. It is a general LLM instead of reasoning LLM like OpenAI-o1 and DeepSeek-R1 (DeepSeek-AI, 2025). $\bullet$ Memory: We retain and flatten all memories in the prompt. $\bullet$ Transition: A scene will transit to the next when all plots in the plot chain have been completed.
@@
Playwriting-Guided 32% 24% 32% 22% 48% 16% 34% 20% 92%
    w/o. critic & revise 24% 24% 26% 34% 18% 28% 12% 32% 66%
@@
Table 3: Win-rates of different dramatic story generation methods over 50 topics. We use the win-rate “B” to represent the ratio that a method performs the best and use the win-rate “W” to represent the probability that a method performs the worst. We highlight the best and worst score using bold and underline for each dimension.
@@ 5.2 Evaluation on Role Agents
Table 4 compares the performance of different architectures. For each architecture, we let each human/agent experience the interactive drama once. This results in 60 playing histories totally. We report the average score of each dimension. The refers to applying one-for-all for the first scene and director-actor for the second and third scene. We first focus on the effect of plot-based reflection. It significantly enhances agency from both dimensions (3.5$\rightarrow$4.0 and 3.3$\rightarrow$4.0). We also notice that plot-based reflection is useful in enhancing character attractiveness (3.5$\rightarrow$3.9) and narrative progression (3.9$\rightarrow$4.3). We conjecture that the reflection encourages characters to display a stronger sense of empathy, which thus contributes to higher attractiveness and decent progression.
The last column of Table 4 suggests the efficiency boost achieved by the . We find that the director-character architecture is advantageous in building character attractiveness because the independent thinking process is more likely to give impressive responses. However, it incurs a lower quality of narrative, especially in progression (4.3$\rightarrow$3.6). This is due to the fact that the communication between director and actor agents inevitably results in information loss.
@@
Table 5 further indicates that the remains decent in the face of two types of players. Interestingly, it even higher scores in character attractiveness and player influence against aggressive agents compared to human players. We conjecture that this is because agents are more eager to interact and the high-quality responses in turn impress the annotators. However, the provocative queries might lead to missing of plots, as shown by the drop in narrative progression.
@@ 6 Conclusion
Our paper adopts a limited scope of playwriting techniques. However, in modern drama, they should not be predetermined; they should evolve and be adapted based on various domains. There is a potential gap between two parts of evaluation. We evaluate dramatic story generation based on 50 manually-crafted premises, while role agents are evaluated based on a high-quality script. We have not found an efficient method to quantify the player’s complete experience across both the creation and engagement of the story. For example, inviting the volunteers to craft the story, play the interactive drama, and eventually make comments can be inefficient and unstable. For our demonstration, we develop visual scenery using the Phaser3 game engine. However, the drama we study in this paper focuses on text-based interactive experiences. Our method is based on prompting GPT-4o, one of the most powerful LLMs, but it still exhibits shortcomings in plot-based reflection and making ideal character decisions. Our future work will focus on developing effective data-driven methods to enhance the specific capabilities. The memory system we use in constructing the LLM agents is out-of-date. We will further improve this aspect.